# Verifiably Following Complex Robot Instructions with Foundation Models

### v2

@@ 5.2 Spatial Grounding Module
Figure 3: [A] Our spatial grounding module leverages a VLM to detect all referent occurrences from prior observations of the environment. [B] We then obtain an initial semantic map with all detected referent instances by backprojecting pixels in segmented referent masks unto the 3D map. [C] We resolve each referent’s spatial comparators with respect to the origin coordinate frame of reference. [D] We finally filter out failing instances to obtain a Referent Semantic Map (RSM) that localizes the exact referent instances described in the instruction.
@@
The spatial grounding module detects and localizes specific instances of objects referenced in a given instruction. To achieve this we extract composable referent descriptors from translated LTL formulae and leverage visual language models [33, 34] to detect and generate segmentation masks of all referent occurrences from the robot’s prior observations of the environment. We then backproject pixels in these masks unto our 3D map to obtain an initial semantic map that localizes all instruction object instances. In the example from Figure 2, all occurrences of green_plush_toy, whiteboard and robot are detected, segmented and backprojected unto the map as shown in Figure 3[a&b].
@@
We filter detected occurrences of referents to obtain the specific object instances described in the given instruction by resolving the 3D spatial comparators captured in each referent’s CRD, recall that CRDs are propositional expressions and can be evaluated as true or false. We define eight (8) spatial comparators (isbetween, isabove, isbelow, isleftof, isrightof, isnextto, isinfrontof, isbehind) to reason about the spatial relationships between referents based on their backprojected 3D positions. Since all backprojected positions are relative to an origin coordinate system, our spatial comparators are resolved from the perspective of this origin position as shown in Figure 3[c]. This type of Relative frame of reference (FoR) when describing spatial relationships between objects, in contrast to an Absolute or Intrinsic FoR, is dominant in English [35], hence a logical choice in our work. For each detected referent, the backprojected 3D position of its center mask pixel is used as its representative position and we resolve a given referent with a spatial description by applying the appropriate spatial comparator to all detected pairs of the desired referent and comparison landmark objects. This filtering process yields a Referent Semantic Map (RSM) that localizes specific object instances described in the original instruction as shown in Figure 3[d]. See appendix section A4 for further details.
@@
VLM Verification: Potential misclassifications from object detector VLMs is the main source of error in this module. We do not address interactively correcting VLM misclassifications as that is out of the scope of this work, however we provide 3D visualizations tools that enable end users to visually inspect and verify that constructed referent semantic maps localized referents correctly.
5.3 Task and Motion Planning Module
@@
Figure 4: [A] A given instruction translated into our LTL syntax $\varphi_{l}$ can be compiled into an equivalent finite-state automaton that captures the temporal constraints of the task. We select a path through this automaton with a strategy that incrementally selects the next progression state from the initial state to the accepting state. We then execute manipulation options and navigation behaviors as required by this high-level task plan. [B] To execute navigation objectives we generate a task progression semantic map (TPSM) that augments the environment with state transition constraints, localizing goal (yellow) and avoidance (red) regions. We convert generated TPSMs into 2D obstacle maps for constraint-aware continuous path planning.
@@
Progressive Motion Planner: Our TAMP algorithm (Alg.1) first compiles an LTL formula with parameterized robot skills into an equivalent finite-state automaton (Figure 4[a]), which can be used to generate a plan that is guaranteed to be correct-by-construction. We select a path that traverses this automaton from its initial state to its accepting state, this is a high-level task plan that interleaves the navigation and manipulation objectives required for satisfying the instruction. We implement a simple strategy that incrementally selects the next progression state from the initial state until the final accepting state is reached, ensuring the generated plan faithfully obeys all temporal subgoal objectives. As shown in Figure 4[a] each progressive automaton state is connected by transition edges that represent the logical expression that must hold for the transition to occur. Our algorithm performs the required low-level behavior to achieve each transition, for manipulation objective subgoals, we execute the appropriate parameterized skill. For navigation objective subgoals, we dynamically generate Task Progression Semantic Maps (TPSM) that localize goal and constraint regions in the environment and perform continuous path planning using the Fast Marching Tree algorithm (FMT) [36].
@@
Task Progression Semantic Maps: A TPSM (Alg.2) augments a 3D scene with the navigation constraints specified by logical state transition expressions, enabling a robot to perform goal-directed and constraint-aware navigation. Segmented regions of interest captured in a TPSM can be arbitrarily large via a nearness threshold (in meters) that defines what it means to be at or close to an object. This parameter can be specified as a high-level parameter or included in the language instruction module’s prompting strategy, such that an LLM infers what value to use based on the given instruction. Goal regions captured in a generated TPSM along with object-centric constraints, such as object shape and size, form the initiation set for manipulation options. Similar to our spatial grounding module, TPSMs are agnostic to temporal logic task representations and can be used out of the box with a variety of planning approaches to localise regions of interest and enable semantic constraint aware motion planning. We primarily evaluate our approach on a ground mobile robot, hence we transform 3D TPSMs into 2D geometric obstacle maps, where constraint regions are also treated as obstacle regions (Figure 4[b]). Our approach however, is robot agnostic and allows for direct planning in 3D TPSMs for appropriate embodiments, such as drones.
6 Experimental Evaluation
@@
Table 1: Performance comparison of one-shot instruction understanding and spatial resolution.
@@
Table 2: Performance comparison of one-shot temporal alignment and plan success rate.
@@
Our evaluations test the hypothesis that our system for translating natural language instructions into LTL expressions and dynamically generating semantic maps enables robots to accurately interpret and execute instructions in large-scale environments without prior training in said environments. We focus on three key questions: (1) Can our language module interpret complex ambiguous natural language instructions? (2) Can our spatial grounding module resolve specific object instances described in instructions? (3) Can our TAMP algorithm generate constraint-satisfying plans?
Figure 5: Our interactive reprompting strategy implemented in the symbol verification node regenerates corrective formulas that improve plan success rates with minimal chat turns.
To answer these questions, we compare LIMP with baseline implementations of an LLM task planner (NLMap-Saycan) [10] and an LLM code-writing planner (Code-as-policies) [12], representing state-of-the-art approaches for language-conditioned open ended robot task execution. Both baselines use the same underlying LLM (gpt-4-0613), prompting structure, and in-context learning examples as our language understanding module. For the code-as-policies planner, in-context examples are converted into language model-generated program (LMP) snippets [12]. To ensure competitive performance, we integrate our CRD syntax, spatial grounding module and low-level robot control into these baselines, allowing them to query object positions, use our FMT* path planner, and execute manipulation options. We also evaluate ablations of our two-stage language understanding module, due to the prominence of language understanding on instruction following. In our complete approach, the first stage samples a conventional LTL formula $\phi_{l}$ from an instruction $l$ by dynamically selecting relevant in-context examples from a standard dataset [14] of natural language and LTL pairs based on cosine similarity. Our first ablation selects in-context examples randomly, and our second ablation skips this stage entirely, directly sampling our LTL syntax with parameterized robot skills $\varphi_{l}$.
Figure 6: [A] Sample generated plan for a multi object-goal navigation task. [B] Sample generated plan for a mobile pick-and-place task.
We perform a large-scale evaluation on 150 instructions across five real-world environments. The taskset includes 24 tasks with fine-grained object descriptions (NLMD) from [10], 25 tasks with complex language (NLMC) from [10], 25 tasks with simple structured phrasing (OKRB) from [11], 37 tasks with complex temporal structures (CT) from [16], and an additional 39 tasks with descriptive spatial constraints and temporal structures (CST). To evaluate instruction understanding, we introduce the following performance metrics: Referent resolution accuracy, Avoidance constraint resolution accuracy, and Spatial relationship resolution accuracy. These metrics use the popular word error rate metric (WER) to measure alignment with the ground-truth. Additionally, we define Temporal alignment accuracy and Planning success rate metrics. A plan is temporally aligned if the sequence of subgoal objectives matches the instructor’s intention, and it is deemed successful if it satisfies all spatial and temporal constraints specified in the instruction. Achieving a high plan success rate requires accurate referent and avoidance constraint resolution, spatial grounding, and temporal alignment. We report the average word error rates for each baseline in Table 1 and the percentage of successful plans and temporally accurate plans in Table 2. In addition to the verification benefits of symbolic planning, LIMP outperforms baselines in most tasksets, especially in complex temporal planning and constraint avoidance. We observe that LLM and code-writing planners generate sequential subgoals effectively, but struggle with temporal constraint adherence. In contrast, our approach ensures each robot step adheres to constraints while achieving subgoals. Additional qualitative results are shown in Figure 6, and robot demonstrations are available on our project website.
7 Limitations and Conclusion
@@
Our study revealed several limitations that merit discussion. LIMP is able to interpret non-finite instructions into LTL formulas however our planner is only able to processes co-safe formulas, which are limited to finite sequences. The accuracy of spatial grounding is dependent on VLM object recognition and depth sensing, with any deficiencies affecting performance. LIMP assumes a consistent environment between mapping and execution, making our approach non-reactive to dynamic changes. Addressing this with editable scene representations is a future goal. Moreover, while our Progressive Motion Planning algorithm is complete, it does not provide any optimality guarantees, future work will focus on enhancing planning efficiency.
Foundation models have the potential to drive the next generation of generally capable robots. Our results suggest leveraging these models for their strengths (LLMs for language, VLMs for vision) in combination with proven approaches for safety, explainability, and verifiable behavior synthesis. This approach for using foundation models in robotics is vital to improve safety and trust in robots that will have to work with and for humans.
Supplementary Materials
Appendix A1 Appendix Summary
@@
Appendix A2 Extended Related Works
A2.1 Foundation Models in Robotics
@@
A2.2 Planning Models in Robotics
@@
Appendix A3 Language Instruction Module
@@
A3.1 Interactive Symbol Verification
Verifying sampled LTL formulas is essential, as such we implement an interactive dialog system that presents users with extracted referent composible referent descriptors (CRDs) in sampled formulas as well as the implied task structure––encoded in the sequence of state-machine transition expressions that must hold to progressively solve the task. We translate the task structure into English statements via a simple deterministic strategy that replaces logical connectives and skill predicates from the formula with equivalent English phrases. Users can verify a formula as correct or provide corrective statements which are used to reprompt the LLM to obtain new formulas. Here is the exact text prompt used for reprompting.
@@
As an illustration, the instruction “Bring the green plush toy to the whiteboard in front of it” yields the interactive Referent and Task Structure Verification dialog shown in Figure A.7.
Figure A.7: Sample interactive referent and task structure verification dialog
Appendix A4 Spatial Grounding Module
@@
Listing 4: Implementation description of 3D spatial comparators
Appendix A5 Task and Motion Planning Module
@@
Pick Skill: Once the robot has executed the near skill and is at the object to be manipulated, we take a picture of the current environment to detect the object using the Owl-Vit model. The robot is guaranteed to be facing the object as the computed path plan uses the backprojected object 3D position to compute yaw angles for the robot. After detecting the object in the picture, we obtain a segmentation mask from detected boundary box using the Segment Anything model, and compute the center pixel of this mask. We feed this center pixel to the Boston dynamics grasping API to compute a motion plan to grasp the object.
@@
Appendix A7 Evaluation and Baseline Details
All computation including planning, loading and running pretrained visual language models was done on a single computer equipped with one NVIDIA GeForce RTX 3090 GPU. We leverage GPT-4-0613 as the underlying LLM for our instruction understanding module and all our baselines. In all experiments we set the LLM temperature to 0, however since deterministic greedy token decoding is not guaranteed with GPT4, we perform multiple (3) queries for each instruction and evaluate on the most recurring response (mode response).
@@
A7.1 NLMap-Saycan Implementation Prompt
@@
Listing 5: Exact prompt to implement NlMap-Saycan LLM planner
A7.2 Code-as-Policies Implementation Prompt
@@
Listing 6: Exact prompt to implement Code-as-policies planner
A7.3 Instruction set
We perform a large-scale evaluation on 150 instructions across five real-world environments. The taskset includes 24 tasks with fine-grained object descriptions (NLMD) from [10], 25 tasks with complex language (NLMC) from [10], 25 tasks with simple structured phrasing (OKRB) from [11], 37 tasks with complex temporal structures (CT) from [16], and an additional 39 tasks with descriptive spatial constraints and temporal structures (CST).
@@
Listing 7: Nlmap Detailed Object Tasks (NLMD)
@@
Listing 8: Nlmap Complex Language Understanding Tasks (NLMC)
@@
Listing 9: Ok-Robot Tasks (OKRB)
@@
Listing 10: Lang2LTL Complex temporal Tasks (CT)
@@
Listing 11: Complex Spatiotemporal Tasks (CST)

### v3

@@ IV-B Spatial Grounding Module
Figure 4: [A] Our spatial grounding module leverages a VLM to detect all referent occurrences from prior observations of the environment. [B] An initial semantic map with all detected referent instances is generated by backprojecting pixels in segmented referent masks unto the 3D map. [C] Each referent’s spatial comparators is resolved with respect to the origin coordinate frame of reference. [D] Failing instances are filtered out to obtain a Referent Semantic Map (RSM) that localizes the exact referent instances described in the instruction.
@@
This module detects and localizes specific object instances referenced in an instruction. From the translated LTL formulas, we extract composable referent descriptors (CRDs) and use vision-language models OWL-ViT [32] and SAM [33] to detect and segment all referent occurrences from the robot’s prior observations of the environment. We backproject pixels in these segmentation masks onto our 3D map, creating an initial semantic map of all instruction object instances. From the example in Figure 3, occurrences of green_plush_toy, whiteboard, and robot are detected, segmented, and backprojected onto the map (Figure 4[a&b]).
@@
To obtain the specific object instances described in the instruction, we resolve the 3D spatial comparators in each referent’s CRD––recall that CRDs are propositional expressions and can be evaluated as true or false. We define eight spatial comparators (isbetween, isabove, isbelow, isleftof, isrightof, isnextto, isinfrontof, isbehind) to reason about spatial relationships based on backprojected 3D positions. Since all backprojected positions are relative to an origin coordinate system, our spatial comparators are resolved from the perspective of this origin position as shown in Figure 4[c]. This type of relative frame of reference (FoR) when describing spatial relationships between objects, in contrast to an absolute or intrinsic FoR, is dominant in English [34], and is a logical choice for our work.
Using the 3D position of each referent’s center mask pixel as its representative position, we resolve a given referent with a spatial description by applying the appropriate spatial comparator to all detected pairs of the desired referent and comparison landmark objects. This filtering process yields a Referent Semantic Map (RSM) that localizes specific object instances described in the instruction as shown in Figure 4[d].
@@
VLM Verification: Potential misclassifications from object detector VLMs is the main source of error in this module. We do not address interactively correcting VLM misclassifications as that is out of the scope of this work, but we provide 3D visualization tools that enable users to visually inspect and verify that constructed referent semantic maps correctly localize referents.
IV-C Task and Motion Planning Module
@@
Figure 5: [A] A given instruction translated into our LTL syntax $\varphi_{l}$ can be compiled into an equivalent finite-state automaton that captures the temporal constraints of the task. A path through this automaton is selected with a strategy that incrementally picks the next progression state from the initial state to the accepting state. The robot then executes the manipulation options and navigation behaviors dictated by this high-level task plan. [B] To execute navigation objectives our approach generates a task progression semantic map (TPSM) that augments the environment with state transition constraints, localizing goal (yellow) and avoidance (red) regions. Generated TPSMs are converted into 2D obstacle maps for constraint-aware continuous path planning.
@@
Progressive Motion Planner (PMP): Our TAMP algorithm compiles the LTL formula with parameterized robot skills into an equivalent finite-state automaton (Figure 5[a]) to generate a verifiably correct task and motion plan. A path from the initial to the accepting state in this automaton is a high-level task plan that interleaves navigation and manipulation objectives required to satisfy the instruction. We select such a path with a simple strategy that incrementally selects the next progression state until the accepting state is reached, ensuring the plan obeys all temporal subgoal objectives. As shown in Figure 5[a], automaton states are connected by transition edges representing the logical expressions required for transitions. For each transition, our algorithm executes the necessary low-level behaviors: for manipulation subgoals, it executes the appropriate parameterized skill; for navigation subgoals, it dynamically generates Task Progression Semantic Maps (TPSMs) to localize goal and constraint regions and performs continuous path planning using the Fast Marching Tree algorithm (FMT[35].
@@
Task Progression Semantic Maps (TPSM): A TPSM augments a 3D scene with navigation constraints specified by logical state transition expressions, enabling goal-directed, constraint-aware navigation. Regions of interest in a TPSM are defined using a nearness threshold specifying proximity to an object. This threshold can be set globally or included in the language instruction module’s prompting strategy, allowing an LLM to infer its value based on the instruction. Like our spatial grounding module, TPSMs are agnostic to temporal logic representations and can be used with various planning approaches for semantic constraint-aware motion planning. We primarily evaluate our approach on a ground mobile robot, hence we transform 3D TPSMs into 2D geometric obstacle maps, where constraint regions are treated as obstacles (Figure 5[b]). However, our approach is robot-agnostic and supports direct planning in 3D TPSMs for appropriate embodiments like drones.
V Evaluation
Our evaluations test the hypothesis that translating natural language instructions into LTL expressions and dynamically generating semantic maps enables robots to accurately interpret and execute instructions in large-scale environments without prior training. We focus on three key questions: (1) Can our language instruction module interpret complex, ambiguous instructions? (2) Can our spatial grounding module resolve specific object instances described in instructions? (3) Can our TAMP algorithm generate constraint-satisfying plans?
To answer these questions, we compare LIMP with two baselines: an LLM task planner (NLMap-Saycan [10]) and an LLM code-writing planner (Code-as-Policies [12]), representing state-of-the-art approaches for language-conditioned, open-ended robot task execution. Both baselines use the same LLM (GPT-4-0613), prompting structure, and in-context learning examples as our language instruction module. In Code-as-Policies, in-context examples are converted into language model-generated program (LMP) snippets [12]. To ensure competitive performance, we integrate our CRD syntax, spatial grounding module, and low-level robot control into these baselines, allowing them to query object positions, use our FMT path planner, and execute manipulation skills.
We also evaluate ablations of our two-stage language instruction module due to its importance in instruction following. In our full approach, the first stage prompts an LLM to generate a conventional LTL formula $\phi_{l}$ from instruction $l$ by dynamically selecting relevant in-context examples from a standard dataset [14] based on cosine similarity. Our first ablation selects in-context examples randomly; and the second ablation skips this stage entirely, directly sampling our LTL syntax with parameterized robot skills $\varphi_{l}$ from $l$.
We conduct a large-scale evaluation across five real-world environments on a diverse task set of 150 instructions from multiple prior works [10, 11, 7]. This task set consists of 24 tasks with fine-grained object descriptions (NLMD), 25 tasks with complex language (NLMC), 25 tasks with simple structured phrasing (OKRB), 37 tasks with complex temporal structures (CT) and 39 tasks with descriptive spatial constraints and temporal structures (CST). Below are examples from each task category illustrating the variety in complexity:
1NLMD: Put the brown multigrain chip bag in the woven basket
2NLMC: I like fruits, can you put something I would like on the yellow sofa for me
3OKRB: Move the soda can to the box
4CT: Visit the purple door elevator, then go to the front desk and then go to the kitchen table, in addition you can never go to the elevator once you have seen the front desk
5CST: I have a white cabinet, a green toy, a bookshelf and a red chair around here somewhere. Take the second item I mentioned from between the first item and the third. Bring it the cabinet but avoid the last item at all costs.
To evaluate instruction understanding, we introduce performance metrics: referent resolution accuracy, avoidance constraint resolution accuracy, and spatial relationship resolution accuracy. These metrics utilize the word error rate (WER), widely used in speech recognition to quantify the difference between a reference and a hypothesis transcription by computing the minimal number of substitutions, deletions, and insertions needed to transform the hypothesis into the reference. WER is calculated as $WER=\frac{S+D+I}{N}$, where $S$ is substitutions, $D$ deletions, $I$ insertions, and $N$ is the total number of words in the reference. In our work:
  • Referent resolution accuracy compares extracted referents in the generated LTL formula to ground truth referents.
  • Avoidance constraint resolution accuracy compares referents to avoid in the LTL formula (denoted by the unary negation operator) against ground truth avoidance referents.
  • Spatial relationship resolution accuracy compares generated Composable Referent Descriptors (CRDs) in the LTL specification with ground truth CRD expressions.
We also define temporal alignment accuracy and planning success rate. A plan is temporally aligned if the sequence of subgoals matches the instructor’s intention, and successful if it satisfies all spatial and temporal constraints specified in the instruction. Achieving a high plan success rate is challenging, requiring accurate referent and avoidance constraint resolution, spatial grounding, and temporal alignment. We report average word error rates for each baseline in Table I and the percentage of successful and temporally accurate plans in Table II.
@@ V Evaluation
TABLE I: Performance comparison of one-shot instruction understanding and spatial resolution.
@@
TABLE II: Performance comparison of one-shot temporal alignment and plan success rate.
@@
VI Discussion
Beyond the verification benefits of symbolic planning, LIMP outperforms baselines in most task sets, notably in complex temporal planning and constraint avoidance. While NLMap-Saycan and Code-as-Policies effectively generate sequential subgoals, they struggle with strict temporal constraints—for example, avoiding a specific referent while approaching another. Our approach ensures each robot step adheres to constraints while achieving subgoals, explaining LIMP’s superior performance on CT and CST tasks. As shown in Table II, LIMP underperforms only against NLMap-Saycan in the NLMC task category. This task set, introduced in the same paper as the baseline [10] (which outperforms LIMP), includes instructions with implicit details such as: “I like fruits, can you put something I would like on the yellow sofa for me.” NLMap-Saycan is better suited to infer and generate plans with possible fruit options, whereas our few-shot LTL translation process is not designed for this.
VII Limitations and Conclusion
@@
Although LIMP is capable of interpreting non-finite instructions into LTL formulas, our planner is currently limited to processing co-safe formulas, which handle only finite sequences. The accuracy of spatial grounding relies on the performance of vision-language models (VLMs) for object recognition meaning any shortcomings in these systems can negatively impact results. Additionally, LIMP assumes a static environment between mapping and execution, making it not responsive to dynamic changes—an area we aim to address with future work on editable scene representations. Our Progressive Motion Planning algorithm is complete but does not guarantee optimality; however, our framework can be used with existing TAMP planners to enhance efficiency.
Foundation models hold significant promise for advancing the next generation of autonomous robots. Our results suggest that combining these models—LLMs for language and VLMs for vision—with established methods for safety, explainability, and verifiable behavior synthesis can lead to more reliable and capable robotic systems.
Acknowledgement
This work was supported by the Office of Naval Research (ONR) under REPRISM MURI N000142412603 and ONR #N00014-22-1-2592, as well as the National Science Foundation (NSF) via grant #1955361. Partial funding was also provided by The Robotics and AI Institute.
Appendix
A1 Appendix Summary
@@ Acknowledgement
A2 Extended Related Works
A2-A Foundation Models in Robotics
@@
A2-B Planning Models in Robotics
@@
A3 Language Instruction Module
@@
A3-A Interactive Symbol Verification
Verifying sampled LTL formulas is essential, as such we implement an interactive dialog system that presents users with extracted referent composible referent descriptors (CRDs) in sampled formulas as well as the implied task structure––encoded in the sequence of state-machine transition expressions that must hold to progressively solve the task. We translate the task structure into English statements via a simple deterministic strategy that replaces logical connectives and skill predicates from the formula with equivalent English phrases. Users can verify a formula as correct or provide corrective statements which are used to reprompt the LLM to obtain new formulas. Below is the exact prompt used for regenerating formulas.
@@
To illustrate, the instruction “Bring the green plush toy to the whiteboard in front of it” yields the interactive Referent and Task Structure Verification dialog below:
1**************************
2Instruction Following
3**************************
4Input_instruction: "Bring the green plush toy to the whiteboard in front of it"
5Sampled LTL formula: F(A & F(B & F(C & FD)))
6 A: near[green_plush_toy]
7 B: pick[green_plush_toy]
8 C: near[whiteboard::isinfrontof(green_plush_toy)]
9 D: release[green_plush_toy, whiteboard::isinfrontof(green_plush_toy)]
10
11***************************
12Referent Verification
13***************************
14I extracted this list of relevant objects based on your instruction:
15 * whiteboard::isinfrontof(green_plush_toy)
16 * green_plush_toy
17Does this match your intention? (y/n)
18
19****************************
20Task Structure Verification
21****************************
22Based on my understanding here is the sequence of subgoal objectives needed to satisfy the task:
23Subgoal_1:
24 Logical Expression: A&!B
25 English translation: I should be near the [green_plush_toy] and not have picked up the [green_plush_toy]
26Subgoal_2:
27 Logical Expression: B&!C
28 English translation: I should have picked up the [green_plush_toy] and not be near the [whiteboard::isinfrontof(green_plush_toy)]
29Subgoal_3:
30 Logical Expression: C&!D
31 English translation: I should be near the [whiteboard::isinfrontof(green_plush_toy)] and not have released the [green_plush_toy] at the [whiteboard::isinfrontof(green_plush_toy)]
32Subgoal_4:
33 Logical Expression: D
34 English translation: I should have released the [green_plush_toy] at the [whiteboard::isinfrontof(green_plush_toy)]
35Does this match your intention? (y/n)
Listing 4: Interactive referent and task structure verification dialog.
A4 Spatial Grounding Module
@@ 33 Logical Expression: D
Listing 5: Implementation description of 3D spatial comparators
A5 Task and Motion Planning Module
@@
Pick Skill: Once the robot has executed the near skill and is at the object to be manipulated, it takes a photo of the current environment to detect the object using the Owl-Vit model. The robot is guaranteed to be facing the object as the computed path plan uses the backprojected object 3D position to compute yaw angles for the robot. After detecting the object in the picture, we obtain a segmentation mask from detected boundary box using the Segment Anything model, and compute the center pixel of this mask. We feed this center pixel to the Boston dynamics grasping API to compute a motion plan to grasp the object.
@@
A7 Evaluation and Baseline Details
All computation including planning, loading and running pretrained visual language models was done on a single computer equipped with one NVIDIA GeForce RTX 3090 GPU. We leverage GPT-4-0613 as the underlying LLM for our instruction understanding module and all our baselines. In all experiments we set the LLM temperature to 0, however since deterministic greedy token decoding is not guaranteed with GPT4, we perform three (3) queries for each instruction and evaluate on the most recurring response (mode response).
@@
We visualize some qualitative results of LIMP from our experiments in Figure A.1. We also highlight results in Figure A.2 that illustrates how our interactive symbol verification and reprompting strategy A3-A improves instruction satisfaction with minimal chat turns for different instruction sets.
Figure A.1: [A] Sample generated plan for a multi object-goal navigation task. [B] Sample generated plan for a mobile pick-and-place task.
Figure A.2: Our interactive reprompting strategy implemented in the symbol verification node regenerates corrective formulas that improve plan success rates with minimal chat turns.
A7-A NLMap-Saycan Implementation Prompt
@@
Listing 6: Exact prompt to implement NlMap-Saycan LLM planner
A7-B Code-as-Policies Implementation Prompt
@@
Listing 7: Exact prompt to implement Code-as-policies planner
A7-C Instruction set
We perform a large-scale evaluation on 150 instructions across five real-world environments. The taskset includes 24 tasks with fine-grained object descriptions (NLMD) from [10], 25 tasks with complex language (NLMC) from [10], 25 tasks with simple structured phrasing (OKRB) from [11], 37 tasks with complex temporal structures (CT) from [7], and an additional 39 tasks we propose that have descriptive spatial constraints and temporal structures (CST).
@@
Listing 8: Nlmap Detailed Object Tasks (NLMD)
@@
Listing 9: Nlmap Complex Language Understanding Tasks (NLMC)
@@
Listing 10: Ok-Robot Tasks (OKRB)
@@
Listing 11: Lang2LTL Complex temporal Tasks (CT)
@@
Listing 12: Complex Spatiotemporal Tasks (CST)